OpenStudy (anonymous):

hint please, integral 8sin^4(x)

OpenStudy (anonymous):

look in the back of the book for "power reduction" formula for sine

OpenStudy (anonymous):

pretty good example worked out here http://www.youtube.com/watch?v=4bP1FuASneY these are all decent simple videos

OpenStudy (anonymous):

\[8sin^{3}(x) sin(x) dx\] \[sin^{3}(x)=\sqrt{1-cos^{2}(x)}(1-cos^{2}(x))\] \[put\] \[u=cos(x)\] \[du=-sin(x)dx\] so \[-8 \sqrt{1-u^{2}}(1-u^{2})du\]

OpenStudy (anonymous):

matter of fact this particular video is identical to your problem, except it doesn't have the 8, (which changes nothing) and has 4x instead of x, which makes yours easier

OpenStudy (anonymous):

myininaya enjoys these i think ... lets see

myininaya (myininaya):

\[8(\sin^2(x))^2=8(\frac{1}{2}(1-\cos(2x)))^2\]

myininaya (myininaya):

\[8 \cdot (\frac{1}{2})^2\int\limits_{}^{}(1-2\cos(2x)+\cos^2(2x)) dx\]

OpenStudy (anonymous):

still going to be faced with \[\cos^2(2x)\] so more work to be done

myininaya (myininaya):

\[2(x-2 \frac{1}{2} \sin(2x)+\int\limits_{}^{}\frac{1}{2}(1+\cos(4x)) dx)\]

myininaya (myininaya):

\[2x-2\sin(2x)+\int\limits_{}^{}(1+\cos(4x)) dx\]

myininaya (myininaya):

\[2x-2\sin(2x)+x+\frac{1}{4}\sin(4x)+C\]

OpenStudy (anonymous):

I think one of the easiest way use by parts \[-\int sin^{3}x cos(x) dx+3/4 \int sin^{2}(2x)dx\]

OpenStudy (anonymous):

\[\int sin^{2}(2x) dx=x/2-sin(4x)/8\] \[\int sin^{3}x cosxdx=sin^{4}(x)/4\]